# Models That Prove Their Own Correctness

### v3

@@ 10             Update
Before we continue with formal analysis of Algorithm 2, let us make a few observations.
Firstly, the parameters are updated (line 11) only when an accepting transcript was generated. This means that the learner can first fully generate the transcript (lines 6-7), and then take backwards passes (line 9) only if the transcript was accepted by $V$. This is useful in practice (e.g. when using neural models) as backwards passes are more computationally expensive than forwards passes.
On the other hand, this means that RLVF requires the parameter initialization $\theta_{0}$ to have Verifiability bounded away from 0, so that accepting transcripts are sampled with sufficient probability. Fortunately, such a Self-Proving base model can be learned using TL. This gives a learning paradigm in which a somewhat-Self-Proving base model is learned with TL (with Verifiability $\delta>0$), and then “amplified” to a fully Self-Proving model using RLVF. This can be seen as an adaptation of the method of Nair et al. [2018] to the setting of Self-Proving models.
Secondly, in comparing Algorithms 1 and 2, we see that the latter (RLVF) does not keep track of the probabilities $\alpha_{s}$. This is because, in RL terms, RLVF is an on-policy algorithm; it generates transcripts using the current learned model, unlike TL that samples them from a distribution whose parameterization is unknown to the learner. Hence, the update step in RLVF is simpler than TL.
We now prove that the update step in RLVF maximizes the Verifiability of $P_{\theta}$; this is analogous to Lemma B.4 for TL. We leave it for future work to use Lemma B.6 to obtain convergence bounds on RLVF (analogous to Theorem B.5). As mentioned in Section 4.2, the gap between the lemma and a full convergence theorem (informally) reduces to the problem of obtaining convergence bounds for Policy Gradient methods, a challenging and active research direction (e.g. Agarwal et al. 2021).
Indeed, the update step that the algorithm takes can be expressed as the random vector over which the expectation is taken (in the right hand side).
Lemma B.6 (RLVF gradient estimation).
Fix an input distribution $\mu$ over $\Sigma^{\ast}$ and a verifier $V$ with round complexity $R$ and answer length $L_{a}$. For any transcript $(x,y,q_{1},\dots,a_{R})$ we let $\mathrm{Acc}_{V}(x,y,q_{1},\dots,a_{R})$ denote the indicator random variable which equals 1 if and only if $V$ accepts the transcript. For any model $P_{\theta}$, denote by $\mathrm{ver}(\theta)$ the verifiability of $P_{\theta}$ with respect to $V$ and $\mu$ (Definition 3.4). Then, for any $\theta$,
@@
where $(q_{r},a_{r})_{r=1}^{R}$ are as sampled in lines 5-6 of Algorithm 2, and $\vec{d}_{s}(\theta)$ is as defined in line 8 therein.
Proof.
@@
Recall the transcript generator of $P_{\theta}$, denoted by $\mathcal{T}^{\theta}_{V}$ (see Lemma B.4). By the definitions of Verifiability in Definition 3.4 and $V(x,y,q_{1},\dots,a_{R})$ in the lemma statement,
@@
We will use Equations (3) through (7) in the proof of Lemma B.4. Up to a change in index notation, these show that, for any $\pi^{\ast}$,
@@
Combining Equations 8 and 9, by linearity of expectation we have that
@@
where in the last equality, the probability is over $(q_{r},a_{r})$ sampled as in Algorithm 2, and it follows from the definition of the transcript generator $\mathcal{T}^{\theta}(x)$. ∎
Appendix C Preliminaries on Stochastic Gradient Ascent
For convenience of the reader, we provide a description of Stochastic Gradient Ascent and quote a theorem on its convergence. We adapt the presentation in Shalev-Shwartz and Ben-David [2014], noting that they present Stochastic Gradient Descent in its more general form for non-differentiable unbounded functions.

### v4

@@ 10      Update
Note that the parameters are updated (line 11) only when an accepting transcript was generated. This means that the learner can first fully generate the transcript (lines 6-7), and then take backwards passes (line 9) only if the transcript was accepted by $V$. This is useful in practice (e.g. when using neural models) as backwards passes are more computationally expensive than forwards passes.
On the other hand, this means that RLVF requires the parameter initialization $\theta_{0}$ to have Verifiability bounded away from 0, so that accepting transcripts are sampled with sufficient probability. Fortunately, such a Self-Proving base model can be learned using TL. This gives a learning paradigm in which a somewhat-Self-Proving base model is learned with TL (with Verifiability $\delta\gg 0$), and then “amplified” to a fully Self-Proving model using RLVF. This can be seen as an adaptation of the method of NairMAZA18 to the setting of Self-Proving models.
When comparing Algorithms˜1 and 2, we see that the latter (RLVF) does not keep track of the probabilities $\alpha_{s}$. This is because, in RL terms, RLVF is an on-policy algorithm; it generates transcripts using the current learned model, unlike TL that samples them from a distribution whose parameterization is unknown to the learner. Hence, the update step in RLVF is simpler than TL.
We show that the update step in RLVF maximizes the Verifiability of $P_{\theta}$.
Lemma 4.3 (RLVF gradient estimation).
Fix an input distribution $\mu$ over $\Sigma^{\ast}$ and a verifier $V$ with round complexity $R$ and answer length $L_{a}$. For any transcript $(x,y,q_{1},\dots,a_{R})$ we let $\mathrm{Acc}_{V}(x,y,q_{1},\dots,a_{R})$ denote the indicator random variable which equals 1 if and only if $V$ accepts the transcript. For any model $P_{\theta}$, denote by $\mathrm{ver}(\theta)$ the verifiability of $P_{\theta}$ with respect to $V$ and $\mu$ (Definition˜3.4). Then, for any $\theta$,
@@
where $(q_{r},a_{r})_{r=1}^{R}$ are as in lines 5-6 of Algorithm˜2, and $\vec{d}_{s}(\theta)$ is as defined in line 8 therein.
Note that, because $\mathrm{Acc}_{V}(\cdot)$ is a 0-1 indicator of whether a transcript was accepted, then the right hand side of the above equation is precisely the direction of the step taken in RLVF (line 9). The proof of Lemma˜4.3 can be found in Section˜A.3.
5 Convergence of Transcript Learning
As an application of Lemma˜4.2, we prove that, under certain conditions, Transcript Learning (TL, Algorithm˜1) is expected to output a Self-Proving model. While the theorem relies on simplifying assumptions such as convexity and Lipschitzness, it offers clean mathematical guarantees that illuminate the core dynamics of Transcript Learning. As we discuss following the theorem statement, such conditions are common in theoretical machine learning literature, allowing us to build intuition even when the assumptions may not hold in practice.
Theorem 5.1 (informal).
Fix a verifier $V$, an input distribution $\mu$, and an autoregressive model family $\{P_{\theta}\}_{\theta\in\mathbb{R}^{d}}$. Fix an honest transcript generator $\mathcal{T}^{\ast}_{V}$. Assume the following:
  1. 1.
    The agreement $A(\theta)$, informally defined as the probability that $P_{\theta}$ generates transcripts agreeing with $\mathcal{T}^{\ast}_{V}$, is concave in $\theta$. Additionally, the logits of $P_{\theta}$ are $B_{\mathrm{Lip}}$-Lipschitz in $\theta$.
  2. 2.
    There exist parameters $\theta^{\ast}$ with $\theta^{\ast}||\leq B_{\mathrm{Norm}}$ such that $P_{\theta^{\ast}}$ is $(1-\varepsilon/2)$-Self Proving.
  3. 3.
    The number of tokens sent by the prover in the proof system is at most $C$.
Then, in expectation, TL run on $O(C^{2}B_{\mathrm{Norm}}^{2}B_{\mathrm{Lip}}^{2}/\varepsilon^{2})$ samples outputs a $(1-\varepsilon)$-Self Proving model.
The full statement and proof of Theorem˜5.1 are deferred to Section˜A.4. Its conditions can be split into two parts. First (item 1), convexity and Lipschitzness, which are simplifying assumptions commonly needed to prove SGD convergence. While convexity does not hold in general for DNNs, analyzing convex settings provides clean mathematical tools for establishing foundational results—an approach commonly used in ML theory, particularly for DNNs. Indeed, several works have addressed the problem of proving convergence without convexity (du2019gradient; bartlett2006convexity; khaled2023better).
Norm-boundedness (item 2), on the other hand, is a (necessary) realizability assumption: if the architecture $\{P_{\theta}\}_{\theta}$ cannot be instantiated with parameters $\theta^{\ast}$, then it cannot be trained to be Self-Proving. This assumption is well-grounded for transformer architectures, as recent theoretical work has established their Turing-completeness (BhattamishraPG20; DehghaniGVUK19).
Finally (item 3), the bound $C$ on the communication complexity of the prover in the Interactive Proof system. This parameter directly affects the efficiency of TL, as reflected in the number of iterations (and sampled transcripts): it depends on both the optimization landscape complexity $B_{\mathrm{Norm}}^{2}B_{\mathrm{Lip}}^{2}/\varepsilon^{2}$ and the communication complexity $C^{2}$. Reducing communication has long been a central objective in the study of proof systems (e.g., GoldreichH98; GoldreichVW02; RRR16). Theorem˜5.1 formalizes how communication-efficient proof systems improve the performance of Self-Proving models.
Remark 5.2 (Towards a convergence theorem for RLVF).
RLVF can be derived by viewing Self-Proving as a reinforcement learning problem in which the agent (prover) is rewarded when the verifier accepts. Indeed, RLVF is the Policy Gradient method (SuttonMSM99) for a verifier-induced reward. Convergence bounds for Policy Gradient methods are a challenging and active area of research (e.g. AgarwalKLM21), and so we leave the full analysis to future work.
6 Learning From Annotations
However, when training Self-Proving model, it may be useful for it to first generate an “annotated” answer $\widetilde{a}$ which is then trimmed down to the actual answer $a$ to be sent to the verifier. We adapt Sections˜3 and 4 to this setting via Annotated Transcripts. The TL and RLVF algorithms naturally extend to annotated transcripts as well. Table˜2 shows that annotations significantly improve performance of TL in practice.
Annotations can be viewed as adding Chain-of-Thought (Wei0SBIXCLZ22). As a concrete example, consider our experiments on computing the GCD. As detailed in Section˜7.2, a proof $\pi$ in this setting is the output of an iterative process—the extended Euclidean algorithm—starting from the input $x$: $x\mapsto\pi_{1}\mapsto\pi_{2}\mapsto\dots\mapsto\pi$. The annotation of the proof $\pi$ consists the first $T$ steps $(\pi_{1},\dots,\pi_{T})$ up to some fixed cutoff $T$. These are prepended to the proof and shown to the model during TL training. At inference time, the model is evaluated only on whether it generated the proof $\pi$ correctly.
We formally capture annotations by introducing a transcript annotator and an answer extractor incorporated into the training and inference stages, respectively. Fix a verifier $V$ in an $R$-round proof system with question length $L_{q}$ and answer length $L_{a}$. An annotation system with annotation length $\widetilde{L_{a}}$ consists of a transcript annotator $A$, and an answer extractor $E$.
In terms of efficiency, think of the annotator as an algorithm of the same computational resources as an honest prover in the system (see Definition˜3.2), and the answer extractor as an extremely simple algorithm (e.g., trim a fixed amount of tokens from the annotation).
To use an annotation system the following changes need to be made:
  • At training time, an input $x$ and transcript $\pi$ is annotated to obtain $\widetilde{\pi}\coloneqq A(x,\pi)$, e.g. before the forwards backwards pass in TL (line 3 in Algorithm˜1).
[…]
7 Experiments
Table 2: Self-Proving transformers computing the GCD. We train a 6.3M parameter GPT to compute the GCD of two integers sampled log-uniformly from $10^{4}]$. Vanilla GPT correctly generates the GCD for almost all inputs, but does not prove correctness to a simple verification algorithm. GPT trained with Transcript Learning (GPT+TL) proves its answer 60.3% of the time; adding Reinforcement Learning from Verifier Feedback (+RLVF) increases this to 78.3%; training with Annotated Transcript Learning (GPT+ATL) gives the highest Verifiability score of 96%.
Learning method Correctness Verifiability
GPT (baseline) 99.8% -
GPT+TL 98.8% 60.3%
GPT+TL+RLVF 98.9% 78.3%
GPT+ATL 98.6% 96.0%
7.1 Setup: Training transformers to predict the GCD
Charton24 empirically studies the power and limitations of learning GCDs with transformers. We follow their setup and two conclusions on settings that make for faster learning: Training from the log-uniform distribution, and choosing a base of representation with many prime factors.
We fix a base of representation $B=210$ and use $\mathbf{x}$ to denote an integer $x$ encoded as a $B$-ary string. [^6] B=210B=210 is chosen following Charton24 to be an integer with many prime factors. For sequences of integers, we write $\mathbf{(x_{1}x_{2})}$ to denote the concatenation of $\mathbf{x_{1}}$ with $\mathbf{x_{2}}$, delimited by a special token. The vocabulary size needed for this representation is $|\Sigma|\approx 210$.
We choose the input distribution $\mu$ to be the log-uniform distribution on $10^{4}]$, and train the transformer on sequences of the form $\mathbf{(x_{1}x_{2}y)}$, where $x_{1},x_{2}\sim\mu$ and $y=GCD(x_{1},x_{2})$. This is a scaling-down of Charton24, to allow single GPU training of Self-Proving transformers. In all of our experiments, we use a GPT model (VaswaniSPUJGKP17) with 6.3M parameters trained on a dataset of 1024K samples in batches of 1024. Full details are deferred to Section˜B.2.
Proving correctness of GCD.
Following Charton24 as a baseline, we find that transformers can correctly compute the GCD with over $99\%$ probability over $(x_{1},x_{2})\sim\mu$. To what extent can they prove their answer? To answer this question, we first devise a natural proof system based on Bézout’s theorem. Its specification and formal guarantees are deferred to Section˜B.1. We denote its verification algorithm by $V$, and highlight some important features of the experimental setup:
  • Completeness: For any $x_{1},x_{2},y\in[10^{4}]$ such that $y=GCD(x_{1},x_{2})$, there exists a proof $\mathbf{\pi}$ such that $V(\mathbf{x_{1}x_{2}y\pi})$ accepts. As detailed in Section˜B.1, the proof $\pi$ consists of a pair of integers who are Bézout coefficients for $x_{1},x_{2}$.
Table˜2 shows our main experimental result, which has the following key takeaways:
  1. 1.
    Transcript Learning (TL) for 100K iterations ($\approx\!$100M samples) results in a Self-Proving transformer that correctly proves 60.3% of its answers.
  2. 2.
  3. 3.
    Most efficient is Annotated Transcript Learning, which yielded a model with 96% Verifiability in 100K iterations. We further investigate their effect next.
7.2 Models generalize beyond annotations
Figure 4: Annotated TL Verifiability as a function of the number of samples NN. Each iteration (X axis) is a batch of 1024 samples from a dataset of $\approx$10M sequences. Every 10k iterations, Verifiability was evaluated on a held-out dataset of 1k inputs. $T$ is the number of steps in Annotated Transcript Learning (Figure˜3), and $T=0$ is non-annotated Transcript Learning. Each $T$ was run with three seeds, with mean depicted by the curve and standard error by the shaded area.
The proof $\mathbf{\pi}$ is annotated by including intermediate steps in its computation. Details are deferred to Section˜B.1; roughly speaking, we observe that the proof $\mathbf{\pi}$ for input $\mathbf{(a,b)}$ is obtained as the last element in a sequence $\mathbf{a,b,\pi_{1},\pi_{2},\dots}$ computed by the Euclidean algorithm. We annotate the proof $\mathbf{\pi}$ by prepending to it the sequence of Euclidean steps $\mathbf{(\pi_{1},\dots,\pi_{T})}$ up to some fixed cutoff $T$.
Figure˜3 shows how $T$ affects the Verifiability of the learned model. As suggested by LeeSLLP23, training the model on more intermediate steps results in better performance; in our case, increasing the number of intermediate steps $T$ yields better Self-Proving models. One might suspect that models only learn to execute the Euclidean algorithm in-context. To rule out this hypothesis, we derive an upper bound on the possible efficacy of such limited models. This bound is based on the Euclidean depth of integers $(x_{1},x_{2})$, which we define as the number of intermediate steps that the Euclidean algorithm makes before terminating on input $(x_{1},x_{2})$. Indeed, a model that only learns to compute (in-context) the arithmetic of the Euclidean algorithm would only be able to prove the correctness of inputs $(x_{1},x_{2})$ whose depth does not exceed the annotation cutoff $T$.
Figure˜3 tells a different story: For each cutoff $T$, we estimate the probability that integers $x_{1},x_{2}\sim\mu$ have Euclidean depth at most $T$ on $10^{5}$ sampled pairs. Larger annotation cutoff $T$ increases Verifiability, but all models exceed their corresponding Euclidean depth bound.
7.3 Base of representation
As mentioned previously, Charton24 concludes that, for a given base of representation $B$, transformers correctly compute the GCD of integers $x_{1},x_{2}$ that are products of primes dividing $B$. Simply put, choosing a base $B$ with many different prime factors yields models with better correctness (accuracy), which suggests why base $B=210=2\cdot 3\cdot 5\cdot 7$ yielded the best results. To test if $B$’s factorization has a similar effect on Verifiability, we train transformers on 68 bases varying the number of prime divisors from $\omega(B)=1$ (i.e., $B$ is a prime power) to $\omega(B)=4$. Figure˜5 shows that $\omega(B)$ correlates not just with correctness (Charton24), but also with Verifiability. Although the finding is statistically significant (no overlapping error margins), the difference is by a few percentages; we attribute this to the smaller (10%) number of samples on which models were trained, relative to other experiments.
Figure 6: RLVF Verifiability as a function of the number of samples NN. Starting from a base model with Verifiability 48% (obtained via Transcript Learning), in each iteration a batch of 2048 inputs are sampled; the model generates a proof for each; the Verifier is used to check which proofs are accepted; then, the model parameters are updated accordingly (see Algorithm˜2). Verifiability was evaluated on a held-out dataset of 1k inputs.
8 Conclusion
Trust between a learned model and its user is fundamental. Interactive Proofs (GoldwasserMR85) provide a general framework for establishing trust via verification algorithms. This work shows that models can be trained to formally prove their outputs within such systems—we call these Self-Proving models. Self-Proving models connect the theory of Interactive Proofs with the goal of Trustworthy ML: they offer formal worst-case soundness guarantees; enabling users to be confident when their models generate correct answers—and detect incorrect answers with high probability.
We support our definition with two general-purpose training methods: Transcript Learning (TL) and Reinforcement Learning from Verifier Feedback (RLVF), whose analyses draw on learning theory, RL, and computational complexity. This work can be extended in several directions: finding conditions for the convergence of RLVF, improving sample complexity bounds for TL, or designing altogether different learning algorithms (e.g., by taking advantage of properties of the verifier).
Limitations.
In our current learning methods, each individual ground-truth capability requires training a separate Self-Proving model. A natural generalization of this approach is to adapt our definition and methods to deal with a single generalist Self-Proving model that proves its correctness to multiple verifiers of different ground-truths.
Moreover, the second strategy presented in Section˜4.1.1—namely, generating accepting transcripts first and then constructing matching inputs—has an inherent limitation: the resulting training distribution is biased by the design of the generator. As a consequence, models may end up learning patterns of the construction process, rather than acquiring generalizable reasoning capabilities. This highlights the importance of using sufficiently diverse generators, and of evaluating model performance on out-of-distribution inputs.
Acknowledgments and Disclosure of Funding
We are grateful to Micah Carroll, Mark Rofin, Avishay Tal, Mojtaba Yaghoobzadeh and anonymous reviewers for their helpful comments. This research was supported by DARPA-TA1 under grant no. HR001119S0076, and by the Simons Collaboration on the Theory of Algorithmic Fairness.
Appendix A Formal Proofs
Section˜A.1 formally specifies the setup in which our results reside. We then prove Lemma˜4.2 in Section˜A.2, Lemma˜4.3 in Section˜A.3, and Theorem˜5.1 in Section˜A.4.
A.1 Specification of the Learning Model
In this section, we fully specify the theoretical framework in which our results reside. We define a learner as an algorithm $\Lambda$ with access to a family of autoregressive models $\{P_{\theta}\}_{\theta}$ and samples from the input distribution $x\sim\mu$. In our setting of Self-Proving models (and in accordance with the Interactive Proofs literature), we give the learner the full specification of the verifier $V$. More formally,
Definition A.1 (Self-Proving model learner).
  • A family of autoregressive models $\{P_{\theta}\}_{\theta\in\mathbb{R}^{d}}$ where $d\in\mathbb{N}$ is the number of parameters in the family. For each $\theta$ and $z\in\Sigma^{\ast}$, the random variable $P_{\theta}(z)$ is determined by the logits $\log p_{\theta}(z)\in\mathbb{R}^{|\Sigma|}$. For any $z\in\Sigma^{\ast}$ and $\sigma\in\Sigma$, the learner $\Lambda$ can compute the gradient of the $\sigma^{\text{th}}$ logit, that is, $\nabla_{\theta}\log\Pr_{\sigma^{\prime}\sim p_{\theta}(z)}[\sigma=\sigma^{\prime}]$. In particular, $\log\Pr_{\sigma^{\prime}\sim p_{\theta}(z)}[\sigma=\sigma^{\prime}]$ is always differentiable in $\theta$.
A.2 Proof of Lemma˜4.2
We let $\mathcal{T}^{\theta}_{V}$ denote the transcript generator induced by the model $P_{\theta}$ when interacting with $V$: for each $x$, $\mathcal{T}^{\theta}_{V}(x)$ is the distribution over transcripts of interactions between $V$ and $P_{\theta}$ on input $x$. We stress that $\pi^{*}\sim\mathcal{T}^{\ast}_{V}(x)$ and $\pi\sim\mathcal{T}^{\theta}_{V}(x)$ are transcripts produced when interacting with the same verifier queries; we can think of the verifier as simultaneously interacting with the honest prover and with the model $P_{\theta}$. [^8] The way it is presented in the algorithm, first the verifier is called by $\mathcal{T}^{\ast}_{V}$ and outputs queries $(q^{*}_{1},\dots q^{*}_{R})$, and then the model is prompted with the verifier queries one a time. This maintains soundness, since a proof system is sound as long as the prover does not know the verifier’s queries in advance. In what follows, we use $\pi^{*}\sim\mathcal{T}^{\ast}_{V}(x)$ and $\pi\sim\mathcal{T}^{\theta}_{V}(x)$ to denote two transcripts that share the same queries. That is, taking $\pi^{\ast}=(y^{\ast},q_{1}^{\ast},a_{1}^{\ast},\dots,q_{R}^{\ast},a_{R}^{\ast})$ to denote an accepting transcript sampled from $\mathcal{T}^{\ast}_{V}(x)$, and $\pi=(y,q^{*}_{1},a_{1},\dots,q^{*}_{R},a_{R})$ to denote a random transcript sampled from $\mathcal{T}_{V}^{\theta}(x)$, we say that $\pi$ and $\pi^{*}$ agree if they agree on the prover answers, namely if:
$(y,a_{1},\dots,a_{R})=(y^{\ast},a_{1}^{\ast},\dots,a_{R}^{\ast}).$
This definition implicitly uses the independence of the verifier and model’s randomness. We now prove that TL correctly estimates the gradient of $A(\theta)$ in its update step.
Proof of Lemma˜4.2.
Throughout this proof, expectations and probabilities will be over the same distributions as in the lemma statement. First, we use the law of total probability together with the autoregressive property of $P_{\theta}$ (Section˜4) to switch from probabilities on transcripts, to products of next-token probabilities. Formally, consider a fixed input $x$, an honest transcript $\pi^{\ast}=(y^{\ast},q_{1}^{\ast},a_{1}^{\ast},\dots,q_{R}^{\ast},a_{R}^{\ast})$, and denote a random transcript sampled from $\mathcal{T}_{V}^{\theta}(x)$ when using the same verifier queries by $\pi=(y,q_{1}^{\ast},a_{1},\dots,q_{R}^{\ast},a_{R})$. For any $r\in[R]$ denote the random variable $\mathcal{T}^{\theta,<r}_{V}\coloneqq\mathcal{T}^{\theta}_{V}(yq^{\ast}_{1}a_{1}\cdots a_{r-1}q^{\ast}_{r})$. Then,
where, as noted above, Equation˜2 uses the independence of the verifier and model’s randomness, Equation˜3 uses the autoregressive property of $P_{\theta}$ (Definition˜A.1), and Equation˜4 is by definition of $\alpha_{s}$ and of $a_{0}$. Next, a basic calculus identity gives
This implicitly assumes that $\Pr_{\pi}\left[\pi=\pi^{*}\right]$ is differentiable in $\theta$; indeed, this follows from Definition˜A.1, where the logits of the model were assumed to by differentiable. Let us focus on the rightmost factor. By Equation˜4,
where the last equality is by definition of $\vec{d}_{s}(\theta)$. Combining Equation˜4 and Equation˜5 gives
By the law of total probability and the linearity of the gradient,
which concludes the proof. ∎
A.3 Proof of Lemma˜4.3
@@ A.1 Specification of the Learning Model
Recall the transcript generator of $P_{\theta}$, denoted by $\mathcal{T}^{\theta}_{V}$ (see Lemma˜4.2). By the definitions of Verifiability in Definition˜3.4 and $V(x,y,q_{1},\dots,a_{R})$ in the lemma statement,
@@
We will use Equations (2) through (6) from the proof of Lemma˜4.2. Up to a change in index notation, these show that, for any $\pi^{\ast}$,
@@
Combining Equations˜7 and 8, by linearity of expectation we have that
@@
where in the last equality, the probability is over $(q_{r},a_{r})$ sampled as in Algorithm˜2, and it follows from the definition of the transcript generator $\mathcal{T}^{\theta}(x)$. ∎
A.4 Proof of Theorem˜5.1
We first restate Theorem˜5.1 in full formality.
Theorem A.2.
Fix a verifier $V$, an input distribution $\mu$, an autoregressive model family $\{P_{\theta}\}_{\theta\in\mathbb{R}^{d}}$, and a norm $||\cdot||$ on $\mathbb{R}^{d}$. Fix an honest transcript generator $\mathcal{T}^{\ast}_{V}$, and assume that the agreement function $A(\theta)\coloneqq\Pr\left[\pi=\pi^{*}\right]$ is concave in $\theta$, where the verifier queries are the same in $\pi^{*}$ and $\pi$, and the probability is over $x\sim\mu$, $\pi_{\theta}\sim\mathcal{T}^{\theta}_{V}(x)$, and $\pi^{*}\sim\mathcal{T}^{\ast}_{V}(x)$. For any $\varepsilon>0$, let $B_{\mathrm{Norm}}$, $B_{\mathrm{Lip}}$ and $C$ be upper-bounds such that the following conditions hold:
  • There exists $\theta^{\ast}\in\mathbb{R}^{d}$ with $\theta^{\ast}||<B_{\mathrm{Norm}}$ such that $A(\theta^{\ast})\geq 1-\varepsilon/2$.
[…]
Denote by $\bar{\theta}$ the output of TL running for $N\geq\left(4\cdot C^{2}\cdot B_{\mathrm{Norm}}^{2}\cdot B_{\mathrm{Lip}}^{2}\right)/\varepsilon^{2}$ iterations and learning rate $\lambda=B_{\mathrm{Norm}}/CB_{\mathrm{Lip}}\sqrt{N}$. Then the expected Verifiability (over the randomness of the samples collected by TL) of $\bar{\theta}$ is at least $1-\varepsilon$. That is, $\operatorname*{\mathbb{E}}_{\bar{\theta}}[\mathrm{ver}_{V,\mu}(\bar{\theta)}]\geq 1-\varepsilon$.
The proof of Theorem˜A.2 goes by reduction to Stochastic Gradient Descent (SGD). Lemma˜4.2 showed that the learner can use its only available tools—sampling honest transcripts, emulating the verifier, and differentiating the logits—to optimize the agreement $A(\theta)$. Since $A(\theta)$ lower bounds the Verifiability of $P_{\theta}$, the former can be used as a surrogate for the latter.
For convenience of the reader, we first provide a description of Stochastic Gradient Ascent (equiv. to SGD) and quote a theorem on its convergence. We adapt the presentation in SSS14, noting that they present Stochastic Gradient Descent in its more general form for non-differentiable unbounded functions. The familiar reader may skip directly to the proof in Section˜A.4.3.
A.4.1 Preliminaries on Stochastic Gradient Ascent